OpenStudy (anonymous):

Multiply x^2/x^2+8 * x+6/x

Parth (parthkohli):

\[\rm {x^2 \over x^2 + 8} \times {x + 6 \over x}\]Hmm...\[\rm {\cancel{x^2}^x \over x^2 + 8}\times {x + 6 \over \cancel{x}}\]

Parth (parthkohli):

You are left with\[\rm {x \over x^2 +8}\times {x + 6 \over 1}\]

OpenStudy (anonymous):

Huh?